# ForgeHLS: A Large-Scale, Open-Source Dataset for High-Level Synthesis

### v2

@@ Abstract
High-Level Synthesis (HLS) plays a crucial role in modern hardware design by transforming high-level code into optimized hardware implementations. However, progress in applying machine learning (ML) to HLS optimization has been hindered by a shortage of sufficiently large and diverse datasets. To bridge this gap, we introduce ForgeHLS, a large-scale, open-source dataset explicitly designed for ML-driven HLS research. ForgeHLS comprises over 400,000 diverse designs generated from 536 kernels covering a broad range of application domains. Each kernel includes systematically automated pragma insertions (loop unrolling, pipelining, array partitioning), combined with extensive design space exploration using Bayesian optimization. Compared to existing datasets, ForgeHLS significantly enhances scale, diversity, and design coverage. We further define and evaluate representative downstream tasks, such as Quality of Result (QoR) prediction and automated pragma exploration, clearly demonstrating ForgeHLS’s utility for developing and improving ML-based HLS optimization methodologies.
Index Terms:
High-Level Synthesis, dataset, Large Language Model.I Introduction
@@
High-Level Synthesis (HLS) is a key technology in modern hardware design, enabling the transformation of high-level code, such as C, C++, or SystemC, into optimized hardware implementations. HLS serves as a bridge between software and hardware design, offering designers the ability to automatically generate hardware descriptions, such as Verilog or VHDL, from high-level algorithmic specifications. This process significantly reduces the time and complexity traditionally required in hardware design, making it a pivotal tool for accelerating the development of custom hardware solutions. A critical aspect of HLS is the use of pragmas. Pragmas are directives that guide hardware optimizations without altering algorithmic code, influencing performance, resource utilization, and power consumption. There are two ways to insert pragmas: by manually adding directive files, and by inserting pragmas directly in the source code.
@@
HLS research has focused on several key downstream tasks, including QoR (Quality of Result) prediction [1, 2, 3, 4], and DSE (Design Space Exploration) [5, 6]. QoR prediction aims to estimate the latency and resource utilization. DSE focuses on systematically exploring different HLS pragma configurations to identify optimal design points that balance performance, resource usage and power consumption. Both tasks rely on large and high-quality datasets to improve the efficiency and effectiveness of HLS tools. However, existing datasets for HLS [7, 8, 9, 10] have been limited in scope and size, often restricted to a small set of applications or a narrow range of hardware configurations. This limitation hinders progress in these critical research domains.
@@
To address these challenges, we introduce ForgeHLS, a large-scale, open-source dataset featuring 429,276 HLS designs generated from 536 diverse C++ kernels. ForgeHLS is 10 times larger than existing HLS datasets [8, 9] and spans a broad spectrum of algorithms across various domains.
@@
Diverse Algorithm Coverage. ForgeHLS collects real-world HLS kernels from established benchmarks as well as official Xilinx collections, such as Vitis HLS examples [11]. We provide a comprehensive collection of all above high-quality benchmarks and have unified them into a consistent format. Moreover, we introduce synthetic code to enhance our dataset. We collect a diverse list of algorithms and generate corresponding HLS kernel implementations. We leverage GPT-4o to generate C++ implementations for each algorithm and perform cross-validation to ensure the quality of the synthetic code. This comprehensive collection approach has enabled ForgeHLS to reach 536 unique kernels, representing an order of magnitude increase over existing HLS datasets.
Automated Pragma Insertion. We developed an automated code analysis workflow that extracts key information from HLS code and explore the full design space by generating all possible combinations of pragma types and factors. For large-scale kernels that cannot be fully explored through exhaustive enumeration, we utilize Bayesian optimization to guide the design space exploration. In previous datasets, the combinations of pragmas are typically defined in a script or a separate file [9, 8, 10], requiring users to generate a large amount of feasible HLS designs for training and testing. This process incurs additional substantial effort and time cost. In contrast, we provide feasible HLS designs with corresponding post-HLS performance metrics, and directly insert pragmas into source code to make generated HLS designs eaiser to use. Our direct pragma insertion allows the code to be easily used for large language model (LLM) and other downstream tasks. This automated pragma insertion workflow will also be open-source, ensuring the dataset’s extensibility to support additional pragma types in the future. By incorporating a comprehensive set of pragmas, we generate 429,276 HLS designs that capture both the functional and structural complexities of HLS code.
@@
Multi-task Evaluation. After constructing our comprehensive dataset, a series of HLS downstream tasks can be investigated and conducted. Specifically, we present two representative tasks: HLS QoR (Quality of results) prediction and HLS pragma DSE. These tasks not only showcase the effectiveness of ForgeHLS but also serve as practical examples for reference in other HLS optimization tasks.
@@
In summary, our contributions offer a valuable foundation for the EDA community, promoting advancements in model performance and accelerating machine learning-driven innovation in HLS. The main contributions are:
@@
We propose an automated workflow for pragma insertion, extensively embedding HLS pragmas into our C++ kernel code.
@@
We define two representative downstream tasks to evaluate our dataset.
II Related Work
II-A HLS Datasets
@@
Numerous efforts have been made to develop datasets and frameworks for HLS research, as listed in Table I and Table II. DB4HLS [7] provided a database of over 100,000 HLS design points collected from MachSuite through exhaustive design space exploration. HLSyn [10] offers 42,000 labeled designs from MachSuite and PolyBench. HLSDataset [9] and HLSFactory [8] extended their datasets on more benchmarks. However, only a subset of their kernels has been open-sourced or includes HLS pragma annotations, and access to some of the post-HLS features is limited or unavailable. Moreover, there are relatively small-scale HLS benchmarks that are commonly favored in mainstream training. These include PolyBench [12], CHStone [13], MachSuite [14], and Rosetta [15]. These benchmarks provide kernel source code, and many datasets are derived from these kernels through DSE to construct their datasets.
Current HLS datasets have several key limitations that hinder their effectiveness for machine learning model training. First, their sources are limited, with most datasets coming from benchmark suites. Second, these datasets lack standardization in format, making it difficult to apply and compare the code uniformly across different sources. Third, they often lack essential post-HLS files, such as ADB (Architecture Description Block) files, which are necessary for CDFGs (constructing control-data flow graphs). Table I compares different methods in multiple dimensions.
II-B HLS Downstream Tasks
@@
HLS research has introduced several downstream tasks, such as QoR (quality of results) prediction and DSE (design space exploration) of HLS pragmas. In the QoR prediction task, prior research leverages machine learning to estimate the performance and area for HLS designs early in the design cycle. These methods enable faster and more efficient optimization without running the time-consuming synthesis flow. Different ML techniques, such as gradient-boosted machine (GBM) [16], XGBoost [17], graph neural networks (GNNs) [2, 18, 1], have been explored to predict post-HLS or post-route QoR at the source code level or the intermediate representation (IR) level. However, their models are trained on small-scale datasets constructed with typical HLS benchmarks, which limits their ability to generalize to new designs with unseen code structures.
@@
In the task of DSE of HLS pragmas, various configurations of HLS pragmas are evaluated and compared to find optimal design parameters with best QoR. Different DSE frameworks are proposed to improve search quality and efficiency, utilizing heuristics [19], conventional methods such as genetic algorithms or simulated annealing [20, 21], or ML-based search [16, 22, 18, 1]. Moreover, recent advancements further support automatic pragma insertion after the completion of DSE [23, 24]. Efficient and high-quality DSE reduces the manual effort required for tuning and optimizing HLS code, reducing the development time. However, these frameworks are often built and evaluated using typical HLS benchmarks like PolyBench, which makes it difficult to assess their effectiveness on more complex designs.
| Source | #K | #D | A#D | A#P | A#T |
| Real-World Code | |||||
| CHStone [13] | 4 | 4,684 | 1171 | 24.0 | 5932.7 |
| MachSuite [14] | 19 | 16,803 | 884.4 | 18.9 | 990.8 |
| PolyBench [12] | 26 | 22,897 | 880.7 | 18.6 | 486.2 |
| Rosetta [15] | 5 | 1,849 | 369.8 | 197.8 | 8985.8 |
| Vitis-Examples [11] | 41 | 22,794 | 556.0 | 8.2 | 673.8 |
| Synthetic Code | |||||
| HLS algorithms | 113 | 93,379 | 826.4 | 8.5 | 285.6 |
| Leetcode algorithms | 22 | 18,780 | 853.6 | 5.6 | 203.2 |
| Operators | 154 | 85,099 | 552.6 | 5.5 | 165.7 |
| RTL chip [25] | 16 | 12,922 | 807.6 | 6.9 | 199.8 |
| RTL ip [25] | 5 | 4,874 | 974.8 | 5.6 | 279.5 |
| RTL module [25] | 131 | 145,195 | 1108.4 | 5.2 | 128.2 |
| Ours in Summary | 536 | 429,276 | 800.9 | 9.4 | 394.4 |
| Related Work | |||||
| DB4HLS | 19 | 124,106 | 6531.8 | 6.5* | 2270.7* |
| HLSyn | 42 | 42,000 | 1000.0 | 7.6 | 629.9 |
| HLSDataset | 34 | 18,876 | 555.1 | - | - |
| HLSFactory | 82 | - | - | - | - |
#K: number of kernels; #D: total number of designs; A#D: average number of designs per kernel; A#P: average number of pragmas per kernel; A#T: average number of tokens per kernel; -: unavailable or not reported; *: data calculated by us.
III Dataset
III-A Dataset Overview
As shown in Table II, ForgeHLS is composed of real-world code from established HLS benchmarks [12, 14, 13, 15, 11], along with synthetic code generated from a variety of algorithms [25, 26]. The kernels represent the C++ source code without pragma, while the designs capture the kernel variations created by embedding different types and factors of HLS pragmas. On average, each kernel in ForgeHLS have 800.9 designs, 9.4 pragmas, and 394.4 tokens in average, indicating the complex code structures and diverse pragma design space within the dataset. An overview of the dataset construction workflow is presented in Figure 1.

### v3

@@ Abstract
High-Level Synthesis (HLS) plays a crucial role in modern hardware design by transforming high-level code into optimized hardware implementations. However, progress in applying machine learning (ML) to HLS optimization has been hindered by a shortage of sufficiently large and diverse datasets. To bridge this gap, we introduce ForgeHLS, a large-scale, open-source dataset explicitly designed for ML-driven HLS research. ForgeHLS comprises over 400k diverse designs generated from 846 kernels covering a broad range of application domains, consuming over 200k CPU hours during dataset construction. Each kernel includes systematically automated pragma insertions (loop unrolling, pipelining, array partitioning), combined with extensive design space exploration using Bayesian optimization. Compared to existing datasets, ForgeHLS significantly enhances scale, diversity, and design coverage. We further define and evaluate representative downstream tasks in Quality of Result (QoR) prediction and automated pragma exploration, clearly demonstrating ForgeHLS utility for developing and improving ML-based HLS optimization methodologies. The dataset and code are public at https://github.com/zedong-peng/ForgeHLS.
1 Introduction
@@
High-Level Synthesis (HLS) is a key technology in modern hardware design, enabling the transformation of high-level code, such as C, C++, or SystemC, into optimized hardware implementations, such as Verilog and VHDL. HLS serves as a bridge between software and hardware design, offering designers the ability to automatically generate hardware descriptions from high-level algorithmic specifications. This process significantly reduces the time and complexity traditionally required in hardware design, making it a pivotal tool for accelerating the development of custom hardware solutions. A critical aspect of HLS is the use of pragmas. Pragmas are directives that guide hardware optimizations without altering algorithmic code, influencing performance, resource utilization, and power consumption.
| Feature |
DB4HLS |
HLSyn |
HLSDataset |
HLSFactory |
ForgeHLS |
| Source | |||||
| CHStone | ✘ | ✘ | ▲ | ▲ | ✔ |
| MachSuite | ✔ | ✔ | ▲ | ✔ | ✔ |
| Polybench | ✘ | ✔ | ✔ | ▲ | ✔ |
| Rosetta | ✘ | ✘ | ▲ | ✔ | ✔ |
| Vitis-Examples | ✘ | ✘ | ✘ | ▲ | ✔ |
| Vitis-Library | ✘ | ✘ | ✘ | ✘ | ✔ |
| Synthetic Code | ✘ | ✘ | ✘ | ✘ | ✔ |
| Pre-HLS | |||||
| Code w/ Pragma | ✔ | ▲ | ✘ | ✘ | ✔ |
| Unified File Structure | ✔ | ✔ | ✘ | ✘ | ✔ |
| Post-HLS | |||||
| Latency | ✔ | ✔ | ▲ | ▲ | ✔ |
| Resources | ✔ | ✔ | ▲ | ▲ | ✔ |
| Graph | ✘ | ▲ | ▲ | ✘ | ✔ |
| Verilog Code | ✘ | ✘ | ✔ | ✔ | ✔ |
@@
HLS research has focused on several key downstream tasks, including QoR (Quality of Result) prediction (Gao et al. 2024; Wu et al. 2022; Lin et al. 2022, 2020), and DSE (Design Space Exploration) (Sohrabizadeh et al. 2021; Kuang et al. 2023). QoR prediction aims to estimate the latency and resource utilization from high-level code. DSE focuses on systematically exploring different HLS pragma configurations to identify optimal design points that balance performance, resource usage and power consumption. Both tasks rely on large and high-quality datasets to improve the efficiency and effectiveness of HLS tools. However, existing datasets for HLS (Ferretti et al. 2021; Abi-Karam et al. 2024; Wei et al. 2023; Bai et al. 2023) have been limited in scope and feature, shown in Table 2.
@@
To address these challenges, we introduce ForgeHLS, a large-scale, open-source dataset featuring 459,850 HLS designs generated from 846 diverse C++ kernels. ForgeHLS is more than 10× larger than existing HLS datasets and spans a broad spectrum of algorithms across various domains. Figure 1 illustrates our comprehensive data construction and evaluation methodology.
@@
Diverse Algorithm Coverage. ForgeHLS collects real-world HLS kernels from established benchmarks as well as official AMD Xilinx collections (Xilinx 2021, 2025), and unifies them into a consistent format. Moreover, we introduce synthetic code to enhance our dataset. We collect a diverse list of algorithms, leverage GPT-4o to generate corresponding HLS kernel implementations and perform validation to ensure the quality of the synthetic code.
Automated Pragma Insertion. We developed an automated code analysis workflow that extracts key information from HLS kernel code and explore the full design space by generating all possible combinations of pragma types and factors. For large-scale kernels that cannot be fully explored through exhaustive enumeration, we utilize Bayesian optimization to guide the design space exploration. By inserting comprehensive pragmas, we obtained 459,850 HLS designs.
| #K | #D | Avg. #P | Avg. #T | #S | |
| Related Work | |||||
| DB4HLS | 19 | 124,106 | 6.5* | 2270.7* | 0.1* |
| HLSyn | 42 | 42,000 | 7.6 | 629.9 | 0.1* |
| HLSDataset | 34 | 18,876 | 22.9* | 1144.5* | 150* |
| HLSFactory | 82* | 3,206* | 16.4* | 520.7* | 2.5* |
| Ours in Summary | |||||
| R-ForgeHLS | 405 | 99,601 | 192.4 | 1298.2 | 1169 |
| S-ForgeHLS | 441 | 360,249 | 4.9 | 165.8 | 1008 |
| ForgeHLS | 846 | 459,850 | 95.3 | 718.6 | 2177 |
| ForgeHLS Subset | |||||
| R-ForgeHLS | |||||
| CHStone | 4 | 4,684 | 24.0 | 5932.7 | 40 |
| MachSuite | 19 | 16,803 | 18.9 | 990.8 | 253 |
| PolyBench | 26 | 22,897 | 18.6 | 486.2 | 412 |
| Rosetta | 5 | 1,849 | 197.8 | 8985.8 | 115 |
| Vitis-Examples | 41 | 22,794 | 8.2 | 673.8 | 65 |
| Vitis-Library | 310 | 30,574 | 244.0 | 1365.9 | 284 |
| S-ForgeHLS | |||||
| HLS algorithms | 113 | 93,379 | 8.5 | 285.6 | 492 |
| Leetcode | 22 | 18,780 | 5.6 | 203.2 | 47 |
| Operators | 154 | 85,099 | 5.5 | 165.7 | 150 |
| RTL chip | 16 | 12,922 | 6.9 | 199.8 | 76 |
| RTL ip | 5 | 4,874 | 5.6 | 279.5 | 66 |
| RTL module | 131 | 145,195 | 5.2 | 128.2 | 177 |
@@
Multi-task Evaluation. After constructing our comprehensive dataset, a series of HLS downstream tasks can be investigated and conducted. Specifically, we present two representative tasks: HLS QoR (Quality of results) prediction and HLS Pragma Exploration.
@@
In summary, our contributions offer a valuable foundation for the ML community, promoting advancements in model performance and accelerating machine learning-driven innovation in HLS. Our main contributions are:
@@
We propose an automated workflow for pragma insertion, directly embedding HLS pragmas into our kernel code.
@@
We define two representative HLS downstream tasks to evaluate HLS dataset.
2 Related Work
2.1 HLS Datasets
@@
Numerous efforts have been made to develop datasets and frameworks for HLS research, as listed in Table 1. There are some classic small-scale HLS benchmarks that are commonly favored in mainstream training, including PolyBench (Pouchet and Yuki 2016), CHStone (Hara et al. 2008), MachSuite (Reagen et al. 2014), and Rosetta (Zhou et al. 2018). These benchmarks provide kernel source code, and many HLS datasets are derived from these kernels through DSE to construct their datasets. These hand-crafted kernels are well-structured and suitable for algorithmic analysis and compiler optimization tasks. However, they cover only a narrow range of application domains. For large-scale HLS dataset, DB4HLS (Ferretti et al. 2021) provided a database of over 100,000 HLS design points collected from MachSuite through exhaustive design space exploration. HLSyn (Bai et al. 2023) offers 42,000 labeled designs from MachSuite and PolyBench. HLSDataset (Wei et al. 2023) and HLSFactory (Abi-Karam et al. 2024) extended their datasets on more benchmarks. However, current HLS datasets face three critical limitations: First, limited availability of HLS codes with pragma, combined with the challenges of manual labeling and expert-driven optimization of pragma coefficients, restricts dataset scalability. Second, inconsistent file structures and missing post-HLS artifacts (e.g., ADB files for CDFG reconstruction) hinder reproducibility and the applicability of these datasets for machine learning tasks. Third, existing datasets often lack comprehensive coverage of diverse application domains and HLS pragma usage.
2.2 HLS Downstream Tasks
@@
HLS research focus on two main downstream tasks: QoR (quality of results) prediction and DSE (design space exploration) of HLS pragmas. In the QoR prediction task, prior research leverages machine learning to estimate the performance and area for HLS designs early in the design cycle, without running the time-consuming high-level synthesis. Different ML techniques, such as gradient-boosted machine (Zhong et al. 2017), XGBoost (Dai et al. 2018), graph neural networks (Wu et al. 2022; Sohrabizadeh et al. 2022; Gao et al. 2024), have been explored to predict post-HLS QoR at the source code level or the intermediate representation graph level.
@@
In the task of DSE, various configurations of HLS pragmas are evaluated and compared to find optimal design parameters with best QoR. Different DSE frameworks are proposed to improve search quality and efficiency, utilizing heuristics (Zhao et al. 2017), conventional methods such as genetic algorithms or simulated annealing (Ferrandi et al. 2008; Schafer, Takenaka, and Wakabayashi 2009), or ML-based search (Zhong et al. 2017; Wu, Xie, and Hao 2021; Sohrabizadeh et al. 2022; Gao et al. 2024; Pouget, Pouchet, and Cong 2024; Pouget, Pouchet, and J.Cong 2024). Efficient and high-quality DSE reduces the manual effort required for tuning and optimizing HLS code, reducing the development time.
3 Dataset
3.1 Dataset Overview